Energy level of Bohr's Atomic Model

1
Expert's answer
2021-11-30T17:45:05-0500

The energy of an electron revolving around a nucleus with a charge of "+Ze"


"E=-\frac{Z^2e^4m}{8h^2\epsilon_0^2}\cdot\frac{1}{n^2}" . Answer